an inclined plane making an angle of 30° with the horizontal placed in a uniform horizontal electric field 200N/C. A body of mass 1kg and a charge 5mC is allowed to slide down from a rest at a height of 1m. if the coefficient friction is 0.2, find the time taken to reach the bottom.

1
Expert's answer
2021-12-15T11:18:56-0500

Given quantities:

"\alpha=30^o"

"E = 200\frac{N}{C}"

"m =1kg"

"q=5mC=5*10^{-3}C"

"h=1m"

"\mu=0.2"

"t \space - \space?"

"ma = Fsin\alpha-\mu Fcos\alpha = F(sin\alpha-\mu cos\alpha)=Eq(sin\alpha-\mu cos\alpha)"

"m\frac{2h}{t^2}= Eq(v\sin\alpha-\mu cos\alpha) \to t = \large\sqrt{\frac{2mh}{Eq(sin\alpha-\mu cos\alpha)}}"

"t = \large\sqrt{\frac{2*1*1}{200*5*10^{-3}*(0.5-0.1*\sqrt3)}}" "=3.26s"